OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 2x − 6 = 22 and arrives at the correct solution? A)Divide both sides by −6 and then add 2. The solution is x = 8. B)Subtract 6 from both sides of the equation and then divide by 2. The solution is x = 7. C)Add 6 to both sides of the equation and then divide by 2. The solution is x = 14. D)Add 6 to both sides of the equation and then divide by 22. The solution is x = −14.

Mehek (mehek14):

what do you think?

Mehek (mehek14):

since 6 is being subtracted, you would do the opposite to get it on the other side what's the opposite of subtraction?

OpenStudy (anonymous):

addition

Mehek (mehek14):

so you would add 6

Mehek (mehek14):

so A and B are out

Mehek (mehek14):

next you have 2x which means 2 times x the opposite of multiplication is?

OpenStudy (anonymous):

division

Mehek (mehek14):

so you would first add 6 and then divide by 2

Mehek (mehek14):

so let's solve it \(2x-6=22\\6+22=28\\28/2=14\)

OpenStudy (anonymous):

oh get it its C

Mehek (mehek14):

yup \(\checkmark\)

OpenStudy (anonymous):

THANK U SOO MUCH

Mehek (mehek14):

anytime ^_^